# Ground States of ${\text{Class}\;\mathcal{S}}$ Theory on ADE Singularities and dual Chern-Simons theory

### v1

@@ Abstract
In radial quantization, the ground states of a gauge theory on ADE singularities $\mathbb{R}^{4}/\Gamma$ are characterized by flat connections that are maps from $\Gamma$ to the gauge group. We study ${\text{Class}\;\mathcal{S}}$ theory of type $\mathfrak{a}_{1}=\mathfrak{su}(2)$ on a Riemann surface of genus $g>1$, without punctures. The fundamental building block of ${\text{Class}\;\mathcal{S}}$ theory is the trifundamental Trinion theory - a low energy limit of two M5 branes compactified on the three-punctured Riemann sphere. We show through the superconformal index, that the supersymmetric Casimir energy of the trifundamental theory imposes a constraint on the set of allowed flat connections, which agrees with the prediction of a duality relating the ground state Hilbert space of ${\text{Class}\;\mathcal{S}}$ on ADE singularities to the Hilbert space of a certain dual Chern-Simons theory whose gauge group is given by the McKay correspondence. The conjecture is shown to hold for $\Gamma=\mathbb{Z}_{k}$, agreeing with the previous results of Benini et al. and Alday et al. Non-abelian generalizations of this duality are analyzed by considering the example of the dicyclic group $\Gamma=\text{Dic}_{2}$, corresponding to gauge group SO$(8)$.
@@ 1 Introduction
There is a duality between the ground states of $G=\text{U}(q)$, ${\mathcal{N}}=4$ supersymmetric Yang Mills theory on $\mathbb{R}^{4}/\Gamma$ and the ground states of level $q$ Chern-Simons Witten:1988hf theory with gauge group $G(\Gamma)$ on $T^{2}$ Vafa:1994tf ; Nakajima:1994nid . Here, $\Gamma$ is a finite subgroup of SU$(2)$, and $G(\Gamma)$ is the gauge group given by the McKay correspondence mckay1981graphs . For SYM in radial quantization, the ground states on the orbifold $(S^{3}/\Gamma)\times S^{1}$ are characterized by nontrivial holonomies of the connection. These holonomies are classified by G-monodromy on the base space or equivalently by group homomorphisms ($\pi_{1}(S^{3}/\Gamma)=\Gamma)\rightarrow G$ up to conjugation. The duality is easy to extend to SU$(q)$ group. Compared with the U$(q)$ case, the theory with SU$(q)$ group has fewer ground states since the image of homomorphisms presenting the flat connections need to be given by a matrix with unit determinant. On the Chern-Simons side, the set of dual ground states correspond to suitable quotients of the weight lattice of $G(\Gamma)$, explained later.
@@
For $\Gamma=\text{Dic}_{2}$, we find a similar phenomenon, with one exception – there is a triplet of weights whose fusion coefficient vanishes and yet the corresponding Casimir energy is lower. In fact, the state appears to correspond to an operator of conformal dimension $\Delta<0$! Because it is an operator that cannot move away from the singularity of $\mathbb{R}^{4}/\Gamma$, it does not necessarily violate unitarity. We conjecture that a similar connection between Casimir energies and Chern-Simons fusion rules holds for ${\text{Class}\;\mathcal{S}}$ theory on other ADE singularities and with other gauge groups for which no simple Lagrangian descriptions exist.
One should mention a similar connection between ${\text{Class}\;\mathcal{S}}$ theory on $\mathbb{R}^{4}/\mathbb{Z}_{k}$ and Chern-Simons theory at level $k$ which stems from the 3D/3D correspondence Dimofte:2011py ; Gukov:2015sna ; Gukov:2016lki and computes partition functions of 6d SCFTs on a product of a lens space $L(k,1)\simeq S^{3}/\mathbb{Z}_{k}$ times a 3-manifold $M_{3}$ (which can be taken in particular to be $\Sigma\times S^{1}$) and leads to a relation between a partition function of a 3D theory $T_{N}[M_{3}]$ [the low energy limit of the $(2,0)$ theory on $M_{3}$] and (complexified) Chern-Simons theory on $M_{3}$.
@@ 2 The duality and the statement of the problem
To apply the VW duality, we look at SU$(2)$ Chern-Simons theory at level $2$. As for SU$(2)$, the simple root is twice the fundamental weight. Only two elements lie in the set (1). In terms of the SU$(2)$ Dynkin label, the states correspond toare $\lambda_{1}=0$ and $\lambda_{1}=2$, and we propose that they match the holonomies (4) with $\lambda_{1}=0$ corresponding to $g_{1}$ and $\lambda_{1}=2$ corresponding to $g_{2}$.
@@
3 Superconformal index of ${\text{Class}\;\mathcal{S}}$
@@
For the Casimir energy computation, we will need to keep the ratios $(\log y)/\beta$, $(\log t)/\beta$, $(\log v)/\beta$ fixed.
@@
Table 3: Summary of operators contributing to index (satisfying $\delta=\Delta-2j_{2}-2R+r=0$) with their charges under symmetry generators. The first part of the table refers to ${\mathcal{N}}=2$ vector multiplet components plus Dirac equation for $\lambda^{1}_{\alpha}$, the middle part gives ${\mathcal{N}}=2$ half-hypermultiplet components and the last one gives contribution from spacetime derivatives.
@@
Consider a field $\Phi_{j_{1},j_{2}}$ with SU$(2)_{1}$ and SU$(2)_{2}$ quantum numbers $j_{1},j_{2}$. It can be extended to a field $\Phi_{j_{1},j_{2}}(x)$ on $S^{3}\times S^{1}$ that satisfies, for all $g\in\Gamma$, the invariance condition ${}^{g}\Phi_{j_{1},j_{2}}({}^{g}x)=\Phi_{j_{1},j_{2}}(x)$, where ${}^{g}x$ denotes the geometrical action of $g$ on $x$ following from the orbifold action of $\Gamma$ on the base space $S^{3}$, and ${}^{g}\Phi$ denotes the combined action of $g$ on spin and gauge, the latter being nontrivial for a nontrivial flat connection. Using the state-operator correspondence, we map these modes to operators at the origin of $\mathbb{R}^{4}$ of the form $(\partial\cdots\partial)\Phi_{j_{1},j_{2}}(0)$, with a single field $\Phi_{j_{1},j_{2}}$ and a finite number (possibly zero) of derivative insertions that are invariant under the combined gauge and spin action of all $g\in\Gamma$. The geometric action of $g$ depends only on $j_{1}$ since $\Gamma$ is embedded only in the SU$(2)_{1}$ factor. From the elements that contribute to the single letter index (see Table 3), the derivatives $\partial_{\pm+}$ and the fermionic fields $\lambda^{1}_{\pm}$ of the vector multiplet are charged under $j_{1}$, and will contribute to a nontrivial (geometric) transformation of the single letter operator $(\partial\cdots\partial)\Phi_{j_{1},j_{2}}(0)$ under $g$. Finally, we have to keep track of the fact that expressions of the form $(\partial\cdots\partial)\lambda$ that vanish by virtue of the Dirac equation $\partial_{\alpha\dot{+}}\lambda^{1\alpha}=0$ should not be counted in the index.
@@ 3.1 Counting holonomy sectors for the Lens space
The derivatives $\partial_{\pm+}$ have a nontrivial $j_{1}$ charges, which are ±$\pm\frac{1}{2}$, respectively. With additional derivative insertions, $\Phi\rightarrow\partial_{++}^{n_{1}}\partial_{-+}^{n_{2}}\Phi$, the quantum number $l$ is increased by $\frac{n_{1}-n_{2}}{2}$.
@@ 3.2 Counting holonomy sectors for Dic2
Note that the holonomies 1-4 are abelian, and act trivially on the vector multiplet, since its fields are all in the adjoint representation. The single letter index can be computed in the same way as for $\Gamma=\mathbb{Z}_{k}$. The new ingredient here is that, for each holonomy sector $g_{m}(r)$ and $g_{m}(s)$, a single letter operator must satisfy two conditions in order to contribute to the index, one from the $r$ generator and one from the $s$ generator. Let $\Phi$ denote a generic BPS field with $j_{1}=0$, transforming in the adjoint representation. The constraints on the descendants $\partial_{++}^{n_{1}}\partial_{-+}^{n_{2}}\Phi$ are
$r\partial_{++}^{n_{1}}\partial_{-+}^{n_{2}}\Phi$ $=\partial_{++}^{n_{1}}\partial_{-+}^{n_{2}}g_{m}(r)\Phi g_{m}(r)^{-1},$ (42)
$s\partial_{++}^{n_{1}}\partial_{-+}^{n_{2}}\Phi$ $=\partial_{++}^{n_{1}}\partial_{-+}^{n_{2}}g_{m}(s)\Phi g_{m}(s)^{-1},$ (43)
@@
$r\partial_{++}^{n_{1}}\partial_{-+}^{n_{2}}\Phi$ $=e^{\pi i(n_{1}-n_{2})/2}\partial_{++}^{n_{1}}\partial_{-+}^{n_{1}}\Phi,$ (44)
$s\partial_{++}^{n_{1}}\partial_{-+}^{n_{2}}\Phi$ $=e^{\pi in_{2}}\partial_{++}^{n_{2}}\partial_{-+}^{n_{1}}\Phi.$ (45)
Equation (45) implies that, generically, a descendant $\partial_{++}^{n_{1}}\partial_{-+}^{n_{2}}\Phi$ by itself cannot satisfy the constraint, so one must consider the most general linear combination of single letter descendants (with $j_{1}=0$):
$\Psi_{(j_{1}=0)}\equiv\sum_{n_{1},n_{2}}C_{n_{1}n_{2}}\partial_{++}^{n_{1}}\partial_{-+}^{n_{2}}\Phi,$ (46)
@@
The nonabelian holonomy number $5$ in Table 4 introduces a new ingredient that, different from the lens space case. Unlike the action of the abelian holonomies 1-4 on the BPS field $\Phi$, which is trivial, the action of holonomy 5 can mix different adjoint components of $\Phi=(\Phi_{0},\Phi_{1},\Phi_{-1})$ (where indices for field components refer to an expansion in a Cartan-Weyl basis):
@@
$\Psi^{\prime}_{(j_{1}=0)}$ $\equiv\sum_{n_{1},n_{2},p}C_{n_{1}n_{2}p}\partial_{++}^{n_{1}}\partial_{-+}^{n_{2}}\Phi_{p},$ (52)
@@
$\Psi_{(j_{1}=\pm 1/2)}$ $\equiv\sum_{n_{1},n_{2},\mu}C_{n_{1}n_{2}\mu}\partial_{++}^{n_{1}}\partial_{-+}^{n_{2}}\lambda_{\mu}.$ (53)
@@
$\Psi^{\prime}_{(j_{1}=\pm 1/2)}$ $\equiv\sum_{n_{1},n_{2},\mu,p}C_{n_{1}n_{2}\mu p}\partial_{++}^{n_{1}}\partial_{-+}^{n_{2}}\lambda_{\mu p}.$ (54)
@@
$\Psi\equiv\sum_{n_{1},n_{2},i,j,k}C_{n_{1}n_{2}ijk}\partial_{++}^{n_{1}}\partial_{-+}^{n_{2}}\Phi_{ijk}.$ (55)
@@
The holonomies $g_{a}(r),g_{a}(s)$ act only on the base field components $\Phi_{ijk}$ in $\Psi$ and do not act on the derivatives $\partial_{\pm+}$. The base field $\Phi_{ijk}$ transforms under the holonomy triplet $abc$ as
@@
The holonomy triplets break down into four distinct sets that each have the same supersymmetric Casimir energy:
  • There are 28 states that have the same energy as the trivial holonomy triplet (111). These states are: (122), (133), (144), (234), (155), (255), (355), (455). In total there are 28 states, since we need to count different orderings [e.g., $(155)$, $(515)$, and $(551)$] as distinct states. These states have the lowest positive supersymmetric Casimir energy $E_{0}=\frac{35}{108}$.

### v2

# Berkeley, CA 94720, USA

@@ Abstract
In radial quantization, the ground states of a gauge theory on ADE singularities $\mathbb{R}^{4}/\Gamma$ are characterized by flat connections that are maps from $\Gamma$ to the gauge group. We study ${\text{Class}\;\mathcal{S}}$ theory of type $\mathfrak{a}_{1}=\mathfrak{su}(2)$ on a Riemann surface of genus $g>1$, without punctures. The fundamental building block of ${\text{Class}\;\mathcal{S}}$ theory is the trifundamental Trinion theory - a low energy limit of two M5 branes compactified on the three-punctured Riemann sphere. We show, through the superconformal index, that the supersymmetric Casimir energy of the trifundamental theory imposes a constraint on the set of allowed flat connections, which agrees with the prediction of a duality relating the ground state Hilbert space of ${\text{Class}\;\mathcal{S}}$ on ADE singularities to the Hilbert space of a certain dual Chern-Simons theory whose gauge group is given by the McKay correspondence. The conjecture is shown to hold for $\Gamma=\mathbb{Z}_{k}$, agreeing with the previous results of Benini et al. and Alday et al. A non-abelian generalization of this duality is analyzed by considering the example of the dicyclic group $\Gamma=\text{Dic}_{2}$, corresponding to Chern-Simons gauge group SO$(8)$.
@@ 1 Introduction
There is a duality Vafa:1994tf ; Nakajima:1994nid between the ground states of $G=\text{U}(q)$, ${\mathcal{N}}=4$ supersymmetric Yang Mills theory on $\mathbb{R}^{4}/\Gamma$ and the ground states of level $q$ Chern-Simons theory Witten:1988hf with gauge group $G(\Gamma)$ on $T^{2}$. Here, $\Gamma$ is a finite subgroup of SU$(2)$, and $G(\Gamma)$ is the gauge group given by the McKay correspondence mckay1981graphs . For SYM in radial quantization, the ground states on the orbifold $(S^{3}/\Gamma)\times S^{1}$ are characterized by nontrivial holonomies of the connection. These holonomies are classified by G-monodromy on the base space or equivalently by group homomorphisms ($\pi_{1}(S^{3}/\Gamma)=\Gamma)\rightarrow G$ up to conjugation. The duality is easy to extend to SU$(q)$ group. Compared with the U$(q)$ case, the theory with SU$(q)$ group has fewer ground states since the image of homomorphisms presenting the flat connections need to be given by a matrix with unit determinant. On the Chern-Simons side, the set of dual ground states correspond to suitable quotients of the weight lattice of $G(\Gamma)$, explained later.
@@
The main new result of this paper is for $\Gamma=\text{Dic}_{2}$, where we verify a similar relation, but with one exception – there is a triplet of weights whose fusion coefficient vanishes and yet the corresponding Casimir energy is lower. In fact, the state corresponds to an operator of conformal dimension $\Delta<0$. Because it is an operator that cannot move away from the singularity of $\mathbb{R}^{4}/\Gamma$, it does not necessarily violate unitarity. We conjecture that a similar connection between Casimir energies and Chern-Simons fusion rules holds for ${\text{Class}\;\mathcal{S}}$ theory on other ADE singularities and with other gauge groups for which no simple Lagrangian descriptions exist. However, the existence of the operator with negative conformal dimension creates complications for genus $g>3$. In those cases flat connections can be constructed for which contributions of positive and negative dimensions from all the trinions combine so that the resulting dimension is zero but there is no obvious corresponding state of the dual Chern-Simons theory. For $g>3$ we can only confirm that a subsector of the $\Delta=0$ space of operators matches the space of states of the dual Chern-Simons theory on the same Riemann Surface.
We note that the connection between ${\text{Class}\;\mathcal{S}}$ theory on $\mathbb{R}^{4}/\mathbb{Z}_{k}$ and Chern-Simons theory at level $k$ has been extended to the famous 3D/3D correspondence Dimofte:2011py ; Gukov:2015sna ; Gukov:2016lki whereby more general partition functions of 6d SCFTs on a product of a lens space $L(k,1)\simeq S^{3}/\mathbb{Z}_{k}$ times a 3-manifold $M_{3}$ (which can be taken in particular to be $\Sigma\times S^{1}$) can be computed. Thus, the partition function of a 3D theory $T_{N}[M_{3}]$ [the low energy limit of the $(2,0)$ theory on $M_{3}$] is related via the 3d/3d correspondence to (complexified) Chern-Simons theory on $M_{3}$. But we are not aware of extensions of the 3d/3d correspondence to orbifolds $S^{3}/\Gamma$ with $\Gamma$ from the D-series.
@@ 2 The duality and the statement of the problem
To apply the VW duality, we look at SU$(2)$ Chern-Simons theory at level $2$. As for SU$(2)$, the simple root is twice the fundamental weight. Only two elements lie in the set (1). In terms of the SU$(2)$ Dynkin label, the states correspond to $\lambda_{1}=0$ and $\lambda_{1}=2$, and we propose that they match the holonomies (4) with $\lambda_{1}=0$ corresponding to $g_{1}$ and $\lambda_{1}=2$ corresponding to $g_{2}$.
@@
3 Casimir energies and the superconformal index
@@
For the Casimir energy computation, we will need to keep the ratios $(\log y)/\beta$, $(\log t)/\beta$, $(\log v)/\beta$ fixed. (In fact, for our purposes setting $v=y=1$ and $t=\exp\beta$ will be sufficient.)
@@
Table 3: Summary of operators contributing to the Index (satisfying $\delta=\Delta-2j_{2}-2R+r=0$) with their charges under symmetry generators. The first part of the table refers to ${\mathcal{N}}=2$ vector multiplet components plus Dirac equation for $\lambda^{1}_{\alpha}$, the middle part gives ${\mathcal{N}}=2$ half-hypermultiplet components and the last one gives contribution from spacetime derivatives.
@@
Consider a field $\Phi_{j_{1},j_{2}}$ with SU$(2)_{1}$ and SU$(2)_{2}$ quantum numbers $j_{1},j_{2}$. It can be extended to a field $\Phi_{j_{1},j_{2}}(x)$ on $S^{3}\times S^{1}$ that satisfies, for all $g\in\Gamma$, the invariance condition ${}^{g}\Phi_{j_{1},j_{2}}({}^{g}x)=\Phi_{j_{1},j_{2}}(x)$, where ${}^{g}x$ denotes the geometrical action of $g$ on $x$ following from the orbifold action of $\Gamma$ on the base space $S^{3}$, and ${}^{g}\Phi$ denotes the combined action of $g$ on spin and gauge, the latter being nontrivial for a nontrivial flat connection. Using the state-operator correspondence, we map these modes to operators at the origin of $\mathbb{R}^{4}$ of the form $(\partial\cdots\partial)\Phi_{j_{1},j_{2}}(0)$, with a single field $\Phi_{j_{1},j_{2}}$ and a finite number (possibly zero) of derivative insertions that are invariant under the combined gauge and spin action of all $g\in\Gamma$. The geometric action of $g$ depends only on $j_{1}$ since $\Gamma$ is embedded only in the SU$(2)_{1}$ factor. From the elements that contribute to the single letter index (see Table 3), the derivatives $\partial_{\pm\dot{+}}$ and the fermionic fields $\lambda^{1}_{\pm}$ of the vector multiplet are charged under $j_{1}$, and will contribute to a nontrivial (geometric) transformation of the single letter operator $(\partial\cdots\partial)\Phi_{j_{1},j_{2}}(0)$ under $g$. Finally, we have to keep track of the fact that expressions of the form $(\partial\cdots\partial)\lambda$ that vanish by virtue of the Dirac equation $\partial_{\alpha\dot{+}}\lambda^{1\alpha}=0$ should not be counted in the index.
@@ 3.1 Counting holonomy sectors for the Lens space
The derivatives $\partial_{\pm\dot{+}}$ have a nontrivial $j_{1}$ charges, which are ±$\pm\frac{1}{2}$, respectively. With additional derivative insertions, $\Phi\rightarrow\partial_{+\dot{+}}^{n_{1}}\partial_{-\dot{+}}^{n_{2}}\Phi$, the quantum number $l$ is increased by $\frac{n_{1}-n_{2}}{2}$.
@@ 3.2 Counting holonomy sectors for Dic2
Note that the holonomies 1-4 are abelian, and they act trivially on the vector multiplet, since its fields are all in the adjoint representation. The single letter index can be computed in the same way as for $\Gamma=\mathbb{Z}_{k}$. The new ingredient here is that, for each holonomy sector $g_{m}(r)$ and $g_{m}(s)$, a single letter operator must satisfy two conditions in order to contribute to the index, one from the $r$ generator and one from the $s$ generator. Let $\Phi$ denote a generic BPS field with $j_{1}=0$, transforming in the adjoint representation. The constraints on the descendants ++˙$\partial_{+\dot{+}}^{n_{1}}\partial_{-\dot{+}}^{n_{2}}\Phi$ are
$r\partial_{+\dot{+}}^{n_{1}}\partial_{-\dot{+}}^{n_{2}}\Phi$ =++˙$=\partial_{+\dot{+}}^{n_{1}}\partial_{-\dot{+}}^{n_{2}}g_{m}(r)\Phi g_{m}(r)^{-1},$ (42)
$s\partial_{+\dot{+}}^{n_{1}}\partial_{-\dot{+}}^{n_{2}}\Phi$ =++˙$=\partial_{+\dot{+}}^{n_{1}}\partial_{-\dot{+}}^{n_{2}}g_{m}(s)\Phi g_{m}(s)^{-1},$ (43)
@@
$r\partial_{+\dot{+}}^{n_{1}}\partial_{-\dot{+}}^{n_{2}}\Phi$ $=e^{\pi i(n_{1}-n_{2})/2}\partial_{+\dot{+}}^{n_{1}}\partial_{-\dot{+}}^{n_{1}}\Phi,$ (44)
$s\partial_{+\dot{+}}^{n_{1}}\partial_{-\dot{+}}^{n_{2}}\Phi$ $=e^{\pi in_{2}}\partial_{+\dot{+}}^{n_{2}}\partial_{-\dot{+}}^{n_{1}}\Phi.$ (45)
Equation (45) implies that, generically, a descendant ++˙$\partial_{+\dot{+}}^{n_{1}}\partial_{-\dot{+}}^{n_{2}}\Phi$ by itself cannot satisfy the constraint, so one must consider the most general linear combination of single letter descendants (with $j_{1}=0$):
$\Psi_{(j_{1}=0)}\equiv\sum_{n_{1},n_{2}}C_{n_{1}n_{2}}\partial_{+\dot{+}}^{n_{1}}\partial_{-\dot{+}}^{n_{2}}\Phi,$ (46)
@@
The nonabelian holonomy number $5$ in Table 4 introduces a new ingredient that is different from the lens space case. Unlike the action of the abelian holonomies 1-4 on the BPS field $\Phi$, which is trivial, the action of holonomy 5 can mix different adjoint components of $\Phi=(\Phi_{0},\Phi_{1},\Phi_{-1})$ (where indices for field components refer to an expansion in a Cartan-Weyl basis):
@@
$\Psi^{\prime}_{(j_{1}=0)}$ $\equiv\sum_{n_{1},n_{2},p}C_{n_{1}n_{2}p}\partial_{+\dot{+}}^{n_{1}}\partial_{-\dot{+}}^{n_{2}}\Phi_{p},$ (52)
@@
$\Psi_{(j_{1}=\pm 1/2)}$ $\equiv\sum_{n_{1},n_{2},\mu}C_{n_{1}n_{2}\mu}\partial_{+\dot{+}}^{n_{1}}\partial_{-\dot{+}}^{n_{2}}\lambda_{\mu}.$ (53)
@@
$\Psi^{\prime}_{(j_{1}=\pm 1/2)}$ $\equiv\sum_{n_{1},n_{2},\mu,p}C_{n_{1}n_{2}\mu p}\partial_{+\dot{+}}^{n_{1}}\partial_{-\dot{+}}^{n_{2}}\lambda_{\mu p}.$ (54)
@@
$\Psi\equiv\sum_{n_{1},n_{2},i,j,k}C_{n_{1}n_{2}ijk}\partial_{+\dot{+}}^{n_{1}}\partial_{-\dot{+}}^{n_{2}}\Phi_{ijk}.$ (55)
@@
The holonomies $g_{a}(r),g_{a}(s)$ act only on the base field components $\Phi_{ijk}$ in $\Psi$ and do not act on the derivatives $\partial_{\pm\dot{+}}$. The base field $\Phi_{ijk}$ transforms under the holonomy triplet $abc$ as
@@
3.2.1 Trinion contribution to the Casimir energy on $S^{3}/\text{Dic}_{2}$
After we decompose a Riemann surface $\Sigma$ into trinion vertices connected by edges to construct a Lagrangian for the ${\text{Class}\;\mathcal{S}}$ theory, as in Figure 1, and after selecting an SU$(2)$ flat connection from the list $1,\dots,5$ for each edge (i.e., for each vector multiplet), the total Casimir energy for the particular configuration of flat connections is the sum of contributions from edges and vertices, computed in Appendix B. It is convenient to split the contribution of each edge into two halves, and add the halves to the contributions of the two vertices attached to the edge. After that, to compute the Casimir energy of configuration of holonomies, we just need to add the contribution of the vertices. Each vertex is labeled by a triplet $(ijk)$, with $1\leq i,j,k\leq 5$ corresponding to the SU$(2)$ flat connections on $S^{3}/\text{Dic}_{2}$ of the edges attached to the vertex. From the calculation of Appendix B we get this modified contribution of each vertex, which we list below. In the list, we group the holonomy triplets into four distinct sets that each have the same supersymmetric Casimir energy:
  • There are 28 triplets that have the same energy as the trivial holonomy triplet (111). These are: (122), (133), (144), (234), (155), (255), (355), (455). In total there are 28 combinations, since we need to count different orderings [e.g., $(155)$, $(515)$, and $(551)$] as distinct combinations. These triplets have the lowest positive supersymmetric Casimir energy $E_{0}=\frac{35}{108}$.